# Judging the Judges: Evaluating Alignment and Vulnerabilities in LLMs-as-Judges

### v5

@@ Appendix E Judge LLM Prompt templates
In Figure 9, we show the prompt template used to guide the judge models during the evaluation process of a 400-question sample from the TriviaQA unfiltered dataset.
@@ Appendix F Metrics for judge models
Percent agreement is simply the ratio of the numbers of times two annotators agree with each other relative to the total number of annotations. This ratio can have values between 00 and $1$. For the binary case, the alignment ratio $\rho$ is given as
@@
Scott’s Pi, (Scott, 1955), measures the alignment of two annotators while also taking into account the possibility of agreement by pure chance. This coefficient usually has values above 00 in most real-world situations. The value of Scott’s Pi is given as
$\pi$ $=\frac{p_{o}-p_{e}}{1-p_{e}}=\frac{\left(\frac{TP+TN}{TP+TN+FP+FN}\right)-\left(\frac{(TP+FP)(TP+FN)+(TN+FN)(TN+FP)}{(TP+TN+FP+FN)^{2}}\right)}{1-\left(\frac{(TP+FP)(TP+FN)+(TN+FN)(TN+FP)}{(TP+TN+FP+FN)^{2}}\right)}$ (5)
Here, $p_{o}$ is the relative observed agreement, and $p_{e}$ is the hypothetical probability of chance agreement.
@@ Appendix G Human Annotation Guidelines
We provide the guidelines used for human evaluation below.
Humn annotation guidelines You will be given a question, a set of reference answers and the answer given by an LLM. Your task is to judge if the answer given by the LLM is correct, as if you were the LLMs teacher grading their exam. An answer should be counted as correct if it is semantically equivalent to (one of the) reference answers. In doing so, please follow the following guidelines: Underspecified answers (e.g. "December" instead of "December 20") should be marked incorrect. Answers that have more information than requested (e.g. "December 20, in Paris" instead of "December 20") should be marked correct, provided the extra information is not incorrect or contrasting the rest of the answer. Answers with unnecessary verbosity but correct answers should be marked correct (E.g. ‘‘Thanks for asking this question! The correct answer is: ..."). If you have trouble judging whether the answer is correct, for instance because you feel you are lacking knowledge required to judge so, please indicate so by marking the answer "maybe correct" or ‘‘maybe incorrect", so that we can further review it.
@@ Appendix H Experiment costs
The costs for the different experiments described in this work belong in three categories – GPU-hours for running open-source models on one or more Nvidia A100 GPUs, OpenAI credits for making API calls to OpenAI models, [^6] Pricing details for OpenAI models are available at https://openai.com/api/pricing/ and human hours for manual annotations of benchmark responses. The estimated costs for the final reported experiments are given in Table 4. In addition to this, previous unreported experiments and trials had an approximate cost of 120 GPU-hours, 100 USD in OpenAI credits, and 50 human hours, bringing the total experimental cost for this work to approximately 200 GPU-hours, USD 125 OpenAI credits, and 75 human annotation hours.
Experiment GPU-hours OpenAI credits Human hours
Main benchmarks 5 2 -
Main evaluations 30 8 10
Human alignment 2 - 9
Error analysis 1.5 - 5
Controlled responses 15 - -
Leniency bias 5 5 -
Guideline bias 10 5 1
Reference bias 5 4 1
Total 73.5 24 26
@@ Appendix I Statistical reliability of Evaluation sample
Due to computational constraints discussed in Appendix A and Appendix H, we limit our evaluation set to randomly sampled 400 questions from TriviaQA (Joshi et al., 2017). In this section, we further take 5 samples of 300 randomly selected questions from the evaluation set and calculate the mean and standard deviation of Scott’s Pi. From Table 5, it can be observed that even on down-sampled sets, the Scott’s $\mathbf{\pi}$ values are similar to Figure 1(b). Standard deviation of all the judge models from the mean Scott’s $\mathbf{\pi}$ is also minimal, barring EM lexical match.
@@
We show the scores assigned by each judge model to each exam-taker model, visualised in Figure 1(a) in Table 6.
Table 6: Judge model score card for every exam-taker model.
@@
Appendix K Exam-taker model base vs chat analysis
@@
Base-Chat pair EM Contains Human GPT-4 Turbo Llama-3 70B
@@
Given the human judgments we have available, we take the opportunity to investigate the performance differences between base and their corresponding chat models. In Table 7, we show the scores assigned by various judge models to four base-chat pairs. According to the default metric EM, the base models outperform the chat models by a large margin. Interestingly, while this difference gets smaller when the answers are judged by humans (second column) or GPT-4 Turbo, there is still a substantial difference for all four pairs, suggesting that the difference is not merely an effect of the increased verbosity of the chat models. Further evidence for that hypothesis is provided by Figure 10(b), in which we can see that while 14% of the errors are shared between the base-chat pairs, almost another 14% of the examples get judged correctly by the base models but not by the chat models, while the opposite happens in only 2.5% of the cases. We consider two alternative hypotheses:
@@
To disentangle these two factors, we manually analyse 400 questions for Llama-2 70B and Llama-2 70B-chat, using our earlier error codes. The results, shown in Figure 10(a), sugest that, at least to some extent, the difference between base and chat models is in fact due to ‘unlearning’ of knowledge: while the number of errors is more or less equal among most categories, there is a stark difference in the incorrect entity category. Substantially more often than the base models, the chat models do answer the question with a semantically plausible but incorrect entity. In Table 8-Table 10, we provide examples of such cases. The results do not show any evidence to support the first hypothesis: the number of errors where the answer cannot be parsed or is just entirely incorrect does not differ between base and chat models.
Figure 10: a) Distribution of incorrect question counts by error codes for the Llama2 70B Base vs Chat exam-taker models evaluated on 400 questions. b) Pie chart showing the percentage of questions categorized by the judgment from Base and Chat models.
Table 8: Knowledge unlearning example 1.
@@
Table 9: Knowledge unlearning example 2
@@
Appendix L Exam-taker model ranking correlation
In Table 11, We use the Spearman Rank correlation coefficient (Spearman, 1904) to assess the rankings of the exam-taker models. To validate these rankings, we randomly select 6 out of 9 exam-taker models across 5 samples, subsequently calculating the mean ($\rho$) and standard deviation ($\sigma$) of the rankings. The results reveal that the contains model exhibits the highest stability and $\rho$ among the rankings, while the majority of judge models achieve a coefficient exceeding 0.7, indicating a strong alignment. Notably, smaller models such as Mistral 7B perform on par with GPT-4 Turbo, highlighting the robustness of smaller models in maintaining rankings.
Table 11: Spearman Rank Correlation Coefficient $\rho$.
Judges $\rho$ $\sigma$
Contains 0.99 0.02
Mistral-7B 0.98 0.03
GPT-4 0.98 0.03
Llama2-13B 0.95 0.18
JudgeLM-7B 0.95 0.05
Llama2-7B 0.94 0.04
Llama3.1-70B 0.94 0.07
Llama3-70B 0.93 0.05
Llama3.1-8B 0.89 0.10
Llama3-8B 0.86 0.07
Llama2-70B 0.84 0.13
Gemma-2B 0.71 0.20
EM 0.67 0.13
Appendix M Too much info confuses judges
In Figure 11-14, we report the guidelines we used for the experiments in § 5.2. The simplest prompt used is Without Guidelines v1 (see Figure 11) where we define a sequential and structured process for the judge model. In Without Guidelines v2 (see Figure 12), we add an additional focus on the overall task and outcome as well. For Guidelines without examples (see Figure 13), we provide the judge models with detailed instructions about the task at hand, along with explicit guidelines on how to evaluate the answers. Additionally, for Guidelines with examples(see Figure 14), we also provide examples to the judge models for further reference.
@@
Appendix N Judge models are sensitive to reference order
We investigate the judges’ sensitivity to reference order by providing the same prompt, question and model response to the judge models, but shuffling the reference order in three different permutations. We compute the consistency score of the model as the percentage of questions for which it gives the same judgment all the 3 times. We observe that the model is more likely to evaluate an answer as correct if the corresponding reference appears early in the list of references (see Figure 15). The smaller judge models sometimes fail to capture all the information in the prompt, and provide judgement based on their own knowledge rather than going by the references (see Figure 16).
@@
As described in § 5.4, for the purpose of the leniency bias experiments, we assume that a judge assigns the correct judgment with a probability of $P_{c}$ and randomly assigns the rest of the samples to be “correct” with a probability $P_{+}$. In this section, we derive the mathematical expressions for $P_{c}$ and $P_{+}$. We assume that in the case of misalignment between the evaluation criteria of guidelines and judge models, the probability of getting an evaluation of ‘‘correct’’ is independent of the actual correctness of the answer (i.e. the judge model effectively flips a coin to give out its judgement). For any given benchmark and judge model, we denote the ground-truth score as $s$, and the true positive and true negative rates as $t_{P}$ and $t_{N}$, respectively, all normalized to be between 00 and $1$.
@@
Similarly, the true negatives, where the exam-taker model response is incorrect, and also correctly identified by the judge model to be incorrect, would also be comprised of two cases: 1) The judge evaluates it correctly according to the given evaluation criteria with a probability of $P_{c}$.2) The judge does not evaluate it according to the given criteria with a probability of $1-P_{c}$, but the evaluation still happens to be correct with a probability of $1-P_{+}$. With the total ratio of the incorrect responses being $1-s$, the true negative rate is therefore given by –
@@
Using Equation 7, we can derive the following:
@@
Substituting the value of $P_{+}$ in Equation 6, we get:
@@
The values of $P_{c}$ and $P_{+}$ can be estimated from observed data using the derived expressions. The estimated probabilities using this method, with human evaluation as the reference, are shown in Figure 17(a).
To validate these derived values, we observe the correlation between the estimated values of $P_{c}$ and Scott’s Pi ($\pi$). As shown in Figure 17(b), we observe that the estimated values of $P_{c}$ are highly correlated to the Scott’s $\mathbf{\pi}$ values for the judge models, with a Pearson correlation coefficient of $0.98$.
Judge model $\pi$ $P_{c}$ $P_{+}$

### v6

@@ Appendix E Judge LLM Prompt templates
In Figure˜8, we show the prompt template used to guide the judge models during the evaluation process of a 400-question sample from the TriviaQA unfiltered dataset.
@@ Appendix F Metrics for judge models
Percent agreement is simply the ratio of the numbers of times two annotators agree with each other relative to the total number of annotations. This ratio can have values between 0 and $1$. For the binary case, the alignment ratio $\rho$ is given as
@@
Scott’s Pi, (Scott, 1955), measures the alignment of two annotators while also taking into account the possibility of agreement by pure chance. This coefficient usually has values above 0 in most real-world situations. The value of Scott’s Pi is given below where $p_{o}$ is the relative observed agreement, and $p_{e}$ is the hypothetical probability of chance agreement.
$\pi=\frac{p_{o}-p_{e}}{1-p_{e}}\\ =\frac{\left(\frac{TP+TN}{TP+TN+FP+FN}\right)-\left(\frac{(TP+FP)(TP+FN)+(TN+FN)(TN+FP)}{(TP+TN+FP+FN)^{2}}\right)}{1-\left(\frac{(TP+FP)(TP+FN)+(TN+FN)(TN+FP)}{(TP+TN+FP+FN)^{2}}\right)}$
@@ Appendix G Human Annotation Guidelines
Human annotation guidelines You will be given a question, a set of reference answers and the answer given by an LLM. Your task is to judge if the answer given by the LLM is correct, as if you were the LLMs teacher grading their exam. An answer should be counted as correct if it is semantically equivalent to (one of the) reference answers. In doing so, please follow the following guidelines: Underspecified answers (e.g. "December" instead of "December 20") should be marked incorrect. Answers that have more information than requested (e.g. "December 20, in Paris" instead of "December 20") should be marked correct, provided the extra information is not incorrect or contrasting the rest of the answer. Answers with unnecessary verbosity but correct answers should be marked correct (E.g. ‘‘Thanks for asking this question! The correct answer is: ..."). If you have trouble judging whether the answer is correct, for instance because you feel you are lacking knowledge required to judge so, please indicate so by marking the answer "maybe correct" or ‘‘maybe incorrect", so that we can further review it.
Preliminary research involved iterative refinement of human annotation guidelines to ensure consistency and reproducibility across annotators with general English semantic knowledge. CS graduate students served as annotators for this experiment. We provide the guidelines used for human evaluation below.
@@ Appendix H Experiment costs
The costs for the different experiments described in this work belong in three categories – GPU-hours for running open-source models on one or more Nvidia A100 GPUs, OpenAI credits for making API calls to OpenAI models, [^5] Pricing details for OpenAI models are available at https://openai.com/api/pricing/ and human hours for manual annotations of benchmark responses. The estimated costs for the final reported experiments are given in Table˜5. In addition to this, previous unreported experiments and trials had an approximate cost of 120 GPU-hours, 100 USD in OpenAI credits, and 50 human hours, bringing the total experimental cost for this work to approximately 200 GPU-hours, USD 125 OpenAI credits, and 75 human annotation hours.
@@ Appendix I Statistical reliability of Evaluation sample
Due to computational constraints discussed in Appendix˜A and Appendix˜H, we limit our evaluation set to randomly sampled 400 questions from TriviaQA (Joshi et al., 2017). In this section, we further take 5 samples of 300 randomly selected questions from the evaluation set and calculate the mean and standard deviation of Scott’s Pi. From Table˜4, it can be observed that even on down-sampled sets, the Scott’s $\mathbf{\pi}$ values are similar to Figure˜1(b). Standard deviation of all the judge models from the mean Scott’s $\mathbf{\pi}$ is also minimal, barring EM lexical match.
@@
We show the scores assigned by each judge model to each exam-taker model, visualised in Figure˜1(a) in Table˜6.
Appendix K Exam-taker model base vs chat analysis
Given the human judgments we have available, we take the opportunity to investigate the performance differences between base and their corresponding chat models. In Table˜7, we show the scores assigned by various judge models to four base-chat pairs. According to the default metric EM, the base models outperform the chat models by a large margin. Interestingly, while this difference gets smaller when the answers are judged by humans (second column) or GPT-4 Turbo, there is still a substantial difference for all four pairs, suggesting that the difference is not merely an effect of the increased verbosity of the chat models. Further evidence for that hypothesis is provided by Figure˜9(b), in which we can see that while 14% of the errors are shared between the base-chat pairs, almost another 14% of the examples get judged correctly by the base models but not by the chat models, while the opposite happens in only 2.5% of the cases.
Experiment GPU-hours OpenAI credits Human hours
Main benchmarks 5 2 -
Main evaluations 30 8 10
Human alignment 2 - 9
Error analysis 1.5 - 5
Controlled responses 15 - -
Leniency bias 5 5 -
Guideline bias 10 5 1
Reference bias 5 4 1
Total 73.5 24 26
@@
Table 6: Judge model score card for every exam-taker model.
@@
Base-Chat pair EM Contains Human
GPT-4
Turbo
Llama-3
70B
@@
We consider two alternative hypotheses:
@@
To disentangle these two factors, we manually analyse 400 questions for Llama-2 70B and Llama-2 70B-chat, using our earlier error codes. The results, shown in Figure˜9(a), sugest that, at least to some extent, the difference between base and chat models is in fact due to ‘unlearning’ of knowledge: while the number of errors is more or less equal among most categories, there is a stark difference in the incorrect entity category. Substantially more often than the base models, the chat models do answer the question with a semantically plausible but incorrect entity. In Table˜9-Table˜11, we provide examples of such cases. The results do not show any evidence to support the first hypothesis: the number of errors where the answer cannot be parsed or is just entirely incorrect does not differ between base and chat models.
Appendix L Exam-taker model ranking correlation
In Table˜8, We use the Spearman Rank correlation coefficient (Spearman, 1904) to assess the rankings of the exam-taker models. To validate these rankings, we randomly select 6 out of 9 exam-taker models across 5 samples, subsequently calculating the mean ($\rho$) and standard deviation ($\sigma$) of the rankings. The results reveal that the contains model exhibits the highest stability and $\rho$ among the rankings, while the majority of judge models achieve a coefficient exceeding 0.7, indicating a strong alignment. Notably, smaller models such as Mistral 7B perform on par with GPT-4 Turbo, highlighting the robustness of smaller models in maintaining rankings.
Judges $\rho$ $\sigma$
Contains 0.99 0.02
Mistral-7B 0.98 0.03
GPT-4 0.98 0.03
Llama2-13B 0.95 0.18
JudgeLM-7B 0.95 0.05
Llama2-7B 0.94 0.04
Llama3.1-70B 0.94 0.07
Llama3-70B 0.93 0.05
Llama3.1-8B 0.89 0.10
Llama3-8B 0.86 0.07
Llama2-70B 0.84 0.13
Gemma-2B 0.71 0.20
EM 0.67 0.13
Table 8: Spearman Rank Correlation Coefficient $\rho$.
Appendix M Too much info confuses judges
In Figure˜10-13, we report the guidelines we used for the experiments in §˜5.2. The simplest prompt used is Without Guidelines v1 (see Figure˜10) where we define a sequential and structured process for the judge model. In Without Guidelines v2 (see Figure˜11), we add an additional focus on the overall task and outcome as well. For Guidelines without examples (see Figure˜12), we provide the judge models with detailed instructions about the task at hand, along with explicit guidelines on how to evaluate the answers. Additionally, for Guidelines with examples(see Figure˜13), we also provide examples to the judge models for further reference.
Figure 9: a) Distribution of incorrect question counts by error codes for Llama2 70B Base vs Chat exam-taker models evaluated on 400 questions. b) Pie chart showing the percentage of questions categorized by the judgment from Base and Chat models.
@@
Table 9: Knowledge unlearning example 1.
@@
Table 11: Knowledge unlearning example 3
@@
Appendix N Judge models are sensitive to reference order
We investigate the judges’ sensitivity to reference order by providing the same prompt, question and model response to the judge models, but shuffling the reference order in three different permutations. We compute the consistency score of the model as the percentage of questions for which it gives the same judgment all the 3 times. We observe that the model is more likely to evaluate an answer as correct if the corresponding reference appears early in the list of references (see Figure˜14). The smaller judge models sometimes fail to capture all the information in the prompt, and provide judgement based on their own knowledge rather than going by the references (see Figure˜15).
@@
As described in §˜5.4, for the purpose of the leniency bias experiments, we assume that a judge assigns the correct judgment with a probability of $P_{c}$ and randomly assigns the rest of the samples to be “correct” with a probability $P_{+}$. In this section, we derive the mathematical expressions for $P_{c}$ and $P_{+}$. We assume that in the case of misalignment between the evaluation criteria of guidelines and judge models, the probability of getting an evaluation of ‘‘correct’’ is independent of the actual correctness of the answer (i.e. the judge model effectively flips a coin to give out its judgement). For any given benchmark and judge model, we denote the ground-truth score as $s$, and the true positive and true negative rates as $t_{P}$ and $t_{N}$, respectively, all normalized to be between 0 and $1$.
@@
Similarly, the true negatives, where the exam-taker model response is incorrect, and also correctly identified by the judge model to be incorrect, would also be comprised of two cases: 1) The judge evaluates it correctly according to the given evaluation criteria with a probability of $P_{c}$.2) The judge does not evaluate it according to the given criteria with a probability of $1-P_{c}$, but the evaluation still happens to be correct with a probability of $1-P_{+}$. With the total ratio of the incorrect responses being 1s1-s, the true negative rate is therefore given by –
@@
Using Equation˜5, we can derive the following.
@@
Substituting the value of $P_{+}$ in Equation˜4, we get:
@@
The values of $P_{c}$ and $P_{+}$ can be estimated from observed data using the derived expressions. The estimated probabilities using this method, with human evaluation as the reference, are shown in Figure˜16(a).
To validate these derived values, we observe the correlation between the estimated values of $P_{c}$ and Scott’s Pi ($\pi$). As shown in Figure˜16(b), we observe that the estimated values of $P_{c}$ are highly correlated to the Scott’s $\mathbf{\pi}$ values for the judge models, with a Pearson correlation coefficient of $0.98$.
Judge model $\pi$ $_{c} P+P_{+}$
@@
Figure 17: Leniency bias and answer consistency. Consistency score, defined as the percentage of questions for which the judge model gives the same judgment for three different answer orders.